A bank offers you two rates. The first is a 5% rate on a 30 year...

  1. A bank offers you two rates. The first is a 5% rate on a 30 year self liquidating mortgage for 75% of the 400,000 value.  The second option is 7.5% for 90% financing also on 30 year amortization.  What is the marginal cost of borrowing.

How would you put the question in a financial calculator what would be N,I/y,Pv,fv,Pmt.
Answer #1

For the First part, you input N = 360 (30 *12 months) , I/y = 0.05/12, pv = 400,000*0.75 = 300,000 , fv =0 and calculate PMT

We get PMT (monthly payment) = $1,610.46

For the second part N = 360, 1/y = 0.075/12, pv = 400000*0.9 = 360,000

We get PMT(monbthly payment) = $2,517.17